# Confidence-aware Self-Semantic Distillation on Knowledge Graph Embedding

### v2

@@ 5.1.1. Datasets
5.1.2. Baselines
5.1.3. Metrics
5.1.4. Implementation Details
We implement all baselines with the proposed CSD strategy in PyTorch. All experiments on four datasets are conducted on one NVIDIA GeForce GTX 4090 GPU(24GB). For general link prediction experiments, we utilize the Adam (Adam, ) optimizer. Batch normalization (BN, ) and dropout (dropout, ) are used to mitigate over-fitting. To investigate the model performance in the low-dimension space, we set the embedding dimensions to 25 and 100. Other detailed parameter settings are approximately identical to their original papers. All results of * are retrieved from the paper of DualDE (DualDE, ) or IterDE (liu2023iterde, ). We trained 500 epochs for each knowledge graph model. For the CSD strategy, we set $L$ to $125$, select $k=50,\gamma=0.25$ for 25 dimensions and $k=200,\gamma=0.1$ for 100 dimensions. We search the value of $\beta$ and $\lambda$ in the interval $(0,1)$. For every sub-training process $\beta$ will be reset to the initial setting value and decay over training epochs to take advantage of the temporal attentiveness. The value of $\beta$ multiply by $(1-ep/EP)$ at the beginning of $ep$-th epoch in this sub-training process, where $EP$ is the total number of epoch for each sub-training process.
5.2. Link Prediction Results (RQ1 and RQ2)
We perform link prediction experiments for all baselines combined with our method. The experimental results of 50 and 100 dimensions are shown in Tables 2 and 5. The best results are in bold. Overall, the baseline model combined with our method outperforms all original baselines on these datasets. Specifically, the LowFER model combined with CSD with 25 dimensions has significant performance gains on the WN18RR dataset with improvements of 279.4% (from 0.0160 to 0.0607), 292.2% (from 0.0282 to 0.1106) and 339.7.5% (from 0.0078 to 0.0343) in MRR, H@10 and H@1, respectively.
Comparing CSD with CS-KD and LWR. Our CSD surpasses both CS-KD and LWR across all backbones and datasets, thereby substantiating CSD’s efficacy. It merits mentioning that employing the comparative method, LWR, on the YAGO3-10 dataset proves challenging due to the necessity of recording the model’s output logits for each sample, resulting in considerable memory consumption. And the comparative methods CS-KD and LWR do not surpass the baseline performance. CS-KD is primarily designed for image classification tasks involving limited classes, encounters challenges in the link prediction task. The large number of entities in this task significantly exceeds the class quantity in image classification tasks, resulting in insufficient samples for CS-KD to acquire class-specific knowledge, thereby leading to performance degradation. LWR’s approach to directly distill knowledge from a weak prediction to the model proves effective in image classification tasks, owed primarily to the dark knowledge contained in the weak prediction distribution. Conversely, in the link prediction task, the dark knowledge of a weak prediction distribution declines due to the immense number of entities involved. Consequently, the detrimental impact of a weak prediction distribution outweighs its benefit, some erroneous prediction distribution for knowledge distillation may cause the student model to undergo performance degradation.
Comparing CSD with SD. Our proposed CSD outperform SD on almost every backbones and datasets. SD, akin to our proposed CSD, distills knowledge of the teacher model obtained through the periodic learning rate strategy. Despite SD yielding satisfactory results on certain models, some instances revealed that the Mean Reciprocal Rank (MRR) performance was inferior to the Baseline. The inclusion of a semantic extraction block facilitates the student model in acquiring semantic knowledge based on the last iteration. This Additional advantage allows CSD to surpass SD in performance across the majority of models and datasets.
@@ 5.2. Link Prediction Results (RQ1 and RQ2)
We executed a comparative analysis of our CSD method with existing and well-established distillation techniques frequently utilized in low-dimensional knowledge graph embedding. These methods include: (1) BKD(firstKD, ): The original distillation method. (2) RKD(park2019relational, ): This distillation technique calculates the distillation loss by evaluating the distance and angle of the embedding vector between the teacher and the student models. (3) TA(mirzadeh2020improved, ): This method chooses an intermediary model between the teacher and the student models for the distillation. (4)DualDE (DualDE, ): This implements a two-stage distillation integrated with the structural distillation technique. (5) IterDE (liu2023iterde, ): This technique utilizes a multi-stage distillation strategy. The experiment results are shown in Table 4.
Comparing CSD with existing distillation strategies. Our technique surpasses almost all current distillation approaches. Notably, our method does not employ the teacher model, yet performance still surpasses all techniques that do utilize a teacher model for distillation.
@@ 5.3. Sensitivity Analyses
We utilises the TuckER model in experiments conducted on FB15k-237 and WN18RR datasets, each of 25 dimensions. The hyperparameters $\beta$ and $\lambda$ are adjusted, incrementally from 0, by intervals of 0.1, up to $1.0$. The remaining settings are carried over from Section 5.1.4. As depicted in Fig. 4, there is an observable trend within the MRR that initially ascends and later descends with an increase in $\beta$ and $\lambda$. The maximum increment in MRR arises when $\beta$ and $\lambda$ rise from 0 to 0.1. Importantly, the MRR remains above the values found at $\beta=0$ or $\lambda=0$ across all other hyperparameter iterations, corroborating the efficacy of our proposed method. Further, the results suggest a dependence of the TuckER model’s performance on both $\beta$ and $\lambda$ hyperparameters. Optimal values of these hyperparameters appear to be reliant on the specific dataset, emphasising the necessity for conscientious tuning to achieve the best performance.
@@ 5.3.3. Ablation Study (RQ5)
To demonstrate the advantages of confidence-aware self-semantic distillation, an ablation study is crucial. We performed these experiments using KGE models with the WN18RR and FB15K-237 dataset, adhering to the detailed parameter settings outlined in Section 5.1.4. The results, as presented in Table 6, signify that the methods without snapshot distillation (CSD(w/o $\mathcal{L}_{so}$)) and iterative self-semantic distillation (CSD(w/o $\mathcal{L}_{se}$)) surpass the performance of the original baselines. We also experimented with the method of directly using embeddings for distillation (CSD (w/o Extractor)). This method also surpass the performance of the baselines. None of these methods surpass the performance of our CSD.
@@ 6. Conclusions
In this paper, we propose a confidence-aware self-semantic distillation strategy called CSD which is model-agnostic and can be combined with any KGE model seamlessly. We design a novel semantic extraction block that extracts knowledge from embeddings and filters reliable knowledge by estimating the confidence of previously learned embeddings. The iteratively incorporated and accumulated iteration-based knowledge enhances the performance of the low-dimensional model, thereby improving the quality of link prediction in Knowledge Graphs (KGs). Experiments verify the effectiveness of our CSD.

### v3

@@ 5.2. Link Prediction Results (RQ1 and RQ2)
We executed a comparative analysis of our CSD method with existing and well-established distillation techniques frequently utilized in low-dimensional knowledge graph embedding. The methods include: (1) BKD(firstKD, ): The first knowledge distillation method. (2) RKD(park2019relational, ): This distillation technique calculates the distillation loss by evaluating the distance and angle of the embedding vector between the teacher and the student models. (3) TA(mirzadeh2020improved, ): This method chooses an intermediary model between the teacher and the student models for the distillation. (4)DualDE (DualDE, ): This implements a two-stage distillation integrated with the structural distillation technique. (5) IterDE (liu2023iterde, ): This technique utilizes a multi-stage distillation strategy. The experiment results are shown in Table 5.
Comparing CSD with existing distillation strategies. CSD surpasses almost all current distillation approaches. Notably, our technique does not employ the teacher model, yet performance still surpasses all techniques that utilize a teacher model for distillation.
@@ 5.3. Sensitivity Analyses
In experiments, we apply the TuckER model to FB15k-237 and WN18RR datasets, both of which are of 25 dimensions. The hyperparameters $\beta$ and $\lambda$ are adjusted, incrementally from 0, by intervals of 0.1, up to $1.0$. The remaining settings are carried over from Section 5.1.4. As depicted in Fig. 4, there is an observable trend within the MRR that initially ascends and later descends with an increase in $\beta$ and $\lambda$. The maximum increment in MRR arises when $\beta$ and $\lambda$ rise from 0 to 0.1. Importantly, the MRR remains above the values found at $\beta=0$ or $\lambda=0$ across all other hyperparameter iterations, corroborating the efficacy of our proposed method. Further, the results suggest a dependence of the TuckER model’s performance on both $\beta$ and $\lambda$ hyperparameters. Optimal values of these hyperparameters appear to be reliant on the specific dataset, emphasizing the necessity for conscientious tuning to achieve the best performance.
@@ 5.3.3. Ablation Study (RQ5)
To demonstrate the advantages of our confidence-aware self-semantic distillation, an ablation study is crucial. We performed these experiments using KGE models with the WN18RR and FB15K-237 dataset, adhering to the detailed parameter settings outlined in Section 5.1.4. The results, as presented in Table 6, signify that the methods without snapshot distillation (CSD(w/o $\mathcal{L}_{so}$)) and iterative self-semantic distillation (CSD(w/o $\mathcal{L}_{se}$)) surpass the performance of the original baselines. We also experimented with the method of directly using embeddings for distillation (CSD (w/o Extractor)). This method also surpasses the performance of the baselines. None of these methods surpass the performance of our CSD.
@@ 6. Conclusions
In this paper, we propose a confidence-aware self-semantic distillation strategy called CSD, which is model-agnostic and can be combined with any KGE model seamlessly. We design a novel semantic extraction block that extracts knowledge from embeddings and filters reliable knowledge by estimating the confidence of previously learned embeddings. The iteratively incorporated and accumulated iteration-based knowledge enhances the performance of the low-dimensional model, thereby improving the quality of link prediction in Knowledge Graphs (KGs). Experiments verify the effectiveness of our CSD.